If cot(0)=5/8. How do I find Sine(0)?
Answer by stanbon(75887)

If cot(0)=5/8. How do I find Sine(0)?
-------
By definition, cot = x/y
----------
Since cot = 5/8, x = 5 and y = 8
-----
Then r = sqrt[5^2+8^2] = sqrt[89]
----
Your Problem::
sin(t) = y/r = 8/sqrt(89)
----------------------------------
Cheers,
Stan H.
--------------------